OpenStudy (anonymous):

(a^2n - a^n - 6) / (a^n + 8) The 2n is in exponent form.

OpenStudy (anonymous):

@AJ01

OpenStudy (anonymous):

The numerator can be factored as a quadratic: \[\frac{a^{2n}-a^n-6}{a^n+8}=\frac{(a^n-3)(a^n+2)}{a^n+8}\] By any chance, did you mean to write \(a^{3n}\) in the denominator?

OpenStudy (anonymous):

There is no a^3n.

OpenStudy (anonymous):

I only ask because you could have factored that as a sum of cubes. What exactly are you supposed to do here?

OpenStudy (anonymous):

Divide?

OpenStudy (anonymous):

Alright, let's try synthetic division. First, let's replace \(a^n\) with something simpler, like \(x\): \[\frac{a^{2n}-a^n-6}{a^n+8}=\frac{x^2-x-6}{x+8}\]  which means the quotient is \[x-9+\frac{78}{x+8}=a^n-9+\frac{78}{a^n+8}\]